# Graph Neural Networks for Virtual Sensing in Complex Systems: Addressing Heterogeneous Temporal Dynamics

### v1

@@ Abstract
Real-time condition monitoring is crucial for the reliable and efficient operation of complex systems. However, relying solely on physical sensors can be limited due to their cost, placement constraints, or inability to directly measure certain critical parameters. Virtual sensing addresses these limitations by leveraging readily available sensor data and system knowledge to estimate inaccessible parameters or infer system states. The increasing complexity of industrial systems necessitates deployments of sensors with diverse modalities to provide a comprehensive understanding of system states. These sensors capture data at varying frequencies to monitor both rapid and slowly varying system dynamics, as well as local and global state evolutions of the systems. However, this leads to heterogeneous temporal dynamics, which, particularly under varying operating conditions, pose a significant challenge for accurate virtual sensing. To address these challenges, we propose a novel Heterogeneous Temporal Graph Neural Network (HTGNN) framework for virtual sensing. HTGNN explicitly models signals from diverse sensors as distinct node types within a graph structure, enabling the capture of complex relationships between sensors. Additionally, HTGNN integrates context from operating conditions, derived from exogenous variables such as control settings and external environmental factors into the model architecture. This integration allows HTGNN to adapt to diverse operating and environmental conditions, ensuring accurate and robust virtual sensing. We evaluate the effectiveness of HTGNN using two newly released, publicly available datasets: a test-rig bearing dataset with diverse load conditions for bearing load prediction and a comprehensive year-long simulated dataset for train-bridge-track interaction, aimed at predicting bridge live loads. Our extensive experiments demonstrate that HTGNN significantly outperforms established baseline methods in both bearing and bridge load prediction tasks, particularly under highly varying operating conditions. These results highlight HTGNN’s potential as a robust and accurate virtual sensing approach for complex systems, paving the way for improved monitoring, predictive maintenance, and enhanced system performance.
@@
Virtual Sensing, Soft Sensing, Heterogeneous Temporal Dynamics, Graph Neural Networks (GNNs), Sensor Networks, Complex Systems, Sensor Fusion, Load Prediction, Structural Health Monitoring (SHM), Time-Series Analysis
@@ 1 Introduction
  • Operating condition-aware dynamics: HTGNN captures unique temporal dynamics of each sensor modality under varying operating conditions, accounting for the influence of exogenous variables.
  • Novel datasets for load virtual sensing: We introduce two publicly available datasets: a real-world bearing dataset and a simulated train-bridge-track interaction dataset.
@@ 2 Related Works
In this section, we review state-of-the-art virtual sensing methods, covering both model-based and data-driven approaches, and explore advancements in graph neural networks for time series modeling and their emerging applications in virtual sensing.
@@
Graph neural networks (GNNs), originating in computational chemistry [35] and social networks [36, 37], have rapidly expanded into the Industrial Internet of Things (IIoT) [38]. Their ability to capture spatial-temporal dependencies within sensor networks [19, 39] makes them particularly effective in anomaly and fault detection. For example, a GNN-based model leveraging graph attention mechanisms has shown promise in identifying anomalies within multivariate time series [40]. Another approach employed sensor embeddings for graph construction and prediction-based anomaly detection [41]. Recently, a GNN model incorporating dynamic edges has further improved fault detection by capturing evolving relationships between time series, while also considering operating conditions [42]. However, a persistent challenge remains in addressing heterogeneous temporal dynamics often present in sensor data. Beyond anomaly detection, the inherent ability of GNNS to model interactions within sensor networks has shown promise in a board range of applications. These include predicting remaining useful life of bearing [43], forecasting air quality  [44], estimating states in hydro power plants [45], and calibrating ozone sensors using low-cost sensor networks  [46].
@@ 3.2 Framework Overview
Figure 1: Architecture of the proposed Heterogeneous Temporal Graph Neural Network (HTGNN) for Load Prediction.
@@ 3.5 Heterogeneous Interaction Modelling
To effectively capture the complex relationships between sensor nodes and explicitly account for the influence of operating conditions on these interactions, our HTGNN model strategically models heterogeneous interactions within the temporal graph. We address two distinct types of interactions: homogeneous interactions among nodes with similar signal characteristics (e.g., low-frequency to low-frequency, high-frequency to high-frequency) and heterogeneous interactions across nodes with different signal characteristics (e.g., low-frequency to high-frequency). This approach to interaction modeling leverages the node dynamics previously extracted in Sec. 3.4 (low-frequency nodes from Eq. 9, high-frequency nodes from Eq. 14). This method ensures that our model comprehensively integrates both intra- and inter-modality dynamics, enhancing the accuracy and relevance of the sensor data analysis while respecting the signal characteristics.
@@ 4.2 Case Study 2: Bridge
In this case study, we evaluate our proposed HTGNN model for bridge health monitoring, specifically focusing on estimating live load on the bridge deck from displacement and acceleration signals. Live load refers to the dynamic forces exerted on the bridge by due to traffic, such as the weight of vehicles passing over it. Accurate and timely estimation of live load is essential for assessing a bridge’s structural integrity, identifying potential overloads or fatigue damage, and making informed decisions about maintenance and repair [51]. This capability is essential for ensuring the safety and longevity of bridge structures.
@@ 5 Evaluation Setup
To provide a comprehensive evaluation of our proposed method, we benchmark its performance against several established approaches for virtual sensing [55], each selected for its relevance to the problem domain. We compare against the Bidirectional LSTM (BiLSTM) due to its widespread use in sequence modeling, and the 1D Convolutional Neural Network (1D-CNN) for its effectiveness in signal processing tasks. The 1D Gated CNN (1D-GCNN) is included to assess the impact of gating mechanisms similar to those in our approach. Finally, we include the Multivariate Time-series Graph Attention Network (MTGAT) as a state-of-the-art homogeneous graph neural network, providing a benchmark for the effectiveness of our heterogeneous graph-based approach. For each model, we added a head that converts from the final hidden dimension of the model to the output, implemented as a three-layered MLP. All models utilized the SiLU (Sigmoid Linear Unit) activation function. For further details on the hyperparameter tuning of the baseline methods and our proposed HTGNN model, please refer to A.3.
@@ 6.1.1 Overall Quantitative Performance
Baseline method performance. The complex, interdependent interactions between temperature and vibration measurements in bearing load prediction pose a challenge for many models. These interactions vary not only with load but also with rotational speed. While higher loads generally decrease vibration magnitudes [59] but increase temperature change rates [60], the influence of load on these change rates can be altered by the rotational speed. At higher speeds, both temperature and vibration magnitudes often exhibit steeper increases or decreases in response to load changes. Comparison models, such as MTGAT, which assume homogeneous relations between sensor nodes, struggle to capture these heterogeneous dynamics. Similarly, sequence-based models like 1D-CNN and BiLSTM, primarily focusing on temporal dependencies within a single sensor modality, also face limitations in modeling the complex relationships between temperature and vibration across varying rotational speeds.
Benefits of heterogeneous interaction modeling. In contrast, HTGNN’s architecture explicitly models these heterogeneous relationships and their dependence on operating conditions. This enables HTGNN to effectively extract information from both temperature and vibration sensors across varying rotational speeds, resulting in improved load prediction. By capturing these complex dependencies, HTGNN consistently outperforms other models across all rotational speeds and load types, as evidenced by the lowest NRMSE values with tight confidence intervals (Tab. 1) and the lowest maximum MAPE values (Tab. 2). The low variance in HTGNN’s performance metrics further underscores its robustness and reliability, making it a promising solution for virtual load sensing in complex industrial systems.
@@ 6.1.2 In-Depth Analysis of Load Predictions
Challenges in load prediction. The right two plots of Fig. 8 illustrate the varying degrees of deviation from true load values observed in the performance of baseline models, particularly for axial loads. The relatively flat frequency spectra observed at low speeds (Figure 7) indicate a lower signal-to-noise ratio, which may hinder the performance of frequency-based models like 1D-CNN and 1D-GCNN that rely on prominent frequency-domain features. While 1D-GCNN’s gating mechanism offers some mitigation compared to 1D-CNN, both models remain susceptible to significant deviations under these conditions. MTGAT, a homogeneous GNN, also exhibits occasional extreme errors. This might be due to its aggregation of inherently heterogeneous features (e.g., temperature and vibration), potentially hindering its ability to extract meaningful representations. Additionally, BiLSTM, as a sequential model, may face challenges in extracting meaningful frequency information from vibration signals, particularly when signal trends are less apparent.
@@
Worst case performance assessment. To further assess the robustness of each model, we examine their worst-case performance under challenging operating conditions. Fig. 9 presents the average MAPE for their two most challenging scenarios across five runs. The error bars represent the 95% confidence intervals. Notably, these categories often involve low axial and radial loads, which inherently lead to higher MAPE values due to the nature of this metric (i.e., dividing by a small actual value). For axial load prediction, 1D-CNN and 1D-GCNN consistently struggle, particularly under low axial loads and high rotational speeds. These conditions often lead to significant fluctuations and variations in vibration signal magnitudes. CNN-based methods may not capture effectively due to their primary focus on the amplitude of frequency components rather than the overall signal trends and changes, which are crucial for accurate load prediction. While 1D-GCNN’s gating mechanism offers some improvement over 1D-CNN by modulating extracted features based on rotational speed, it does not consistently rectify this limitation across all operating conditions. Similarly, BiLSTM exhibits high axial MAPE under both high radial loads with high speeds and low radial loads with low speeds. These conditions lead to extreme vibration amplitudes (high or low), which BiLSTM, as a sequential model, may struggle to handle effectively due to the complexities in extracting relevant features from such signals. These findings highlight the complex nonlinear relationship between vibration, temperature, rotational speed, and load, posing a particular challenge for baseline models that lack the ability to effectively capture both high-frequency and low-frequency features simultaneously.
@@ 6.2.1 Overall Performance
Tab. 3 presents the model performance in predicting train loads across different temperature categories using Normalized Root Mean Squared Error (NRMSE) with 95% confidence intervals. Notably, all models exhibit a slight increase in NRMSE as temperature decreases. This could be attributed to the inherent difficulty of modeling bridge behavior at lower temperatures or to the fact that the training dataset was not sufficiently representative and did not contain enough observations of bridge behavior under such conditions. HTGNN consistently demonstrates competitive performance across all temperature ranges, showing particular strength at lower temperatures where other models experience higher NRMSE. 1D-GCNN, which incorporates a gating mechanism, also exhibits strong performance, particularly at higher temperatures. 1D-CNN and BiLSTM generally perform well, except at low temperatures. MTGAT consistently shows the highest NRMSE values. This underperformance, in contrast to its competitive results on the bearing dataset, may be attributed to the increased heterogeneity of the bridge sensor dataset. Unlike the bearing dataset, which is dominated by low-frequency temperature signals, the bridge dataset contains a balanced distribution of high-frequency (acceleration) and low-frequency (displacement) signals. The increased heterogeneity in the bridge sensor data could pose challenges for MTGAT’s homogeneous graph structure, potentially limiting its ability to fully capture the complex interactions between different sensor types.

### v2

@@ Abstract
Real-time condition monitoring is crucial for the reliable and efficient operation of complex systems. However, relying solely on physical sensors can be limited due to their cost, placement constraints, or inability to directly measure certain critical parameters. Virtual sensing addresses these limitations by leveraging readily available sensor data and system knowledge to estimate inaccessible parameters or infer system states. The increasing complexity of industrial systems necessitates deployments of sensors with diverse modalities to provide a comprehensive understanding of system states. These sensors capture data at varying frequencies to monitor both rapid and slowly varying system dynamics, as well as local and global state evolutions of the systems. However, this leads to heterogeneous temporal dynamics, which, particularly under varying operating conditions, pose a significant challenge for accurate virtual sensing. To address these challenges, we propose a novel Heterogeneous Temporal Graph Neural Network (HTGNN) framework for virtual sensing. HTGNN explicitly models signals from diverse sensors as distinct node types within a graph structure, enabling the capture of complex relationships between sensors. Additionally, HTGNN integrates context from operating conditions, derived from exogenous variables such as control settings and external environmental factors into the model architecture. This integration allows HTGNN to adapt to diverse operating and environmental conditions, ensuring accurate and robust virtual sensing. We evaluate the effectiveness of HTGNN using two newly released, publicly available datasets: a test-rig bearing dataset with diverse load conditions for bearing load prediction and a comprehensive year-long simulated dataset for train-bridge-track interaction, aimed at predicting bridge live loads. Our extensive experiments demonstrate that HTGNN significantly outperforms established baseline methods in both bearing and bridge load prediction tasks, particularly under highly varying operating conditions. These results highlight HTGNN’s potential as a robust and accurate virtual sensing approach for complex systems, paving the way for improved monitoring, predictive maintenance, and enhanced system performance. Our code and data are available under https://github.com/EPFL-IMOS/htgnn.
@@
Virtual Sensing, Soft Sensing, Heterogeneous Temporal Dynamics, Graph Neural Networks (GNNs), Exogenous Variable, Sensor Networks, Complex Systems, Multi-Modal Sensor Fusion, Load Prediction, Structural Health Monitoring (SHM), Time-Series Analysis
@@ 1 Introduction
  • Operating condition-aware dynamics: We introduce new general purpose modules to capture the unique temporal dynamics of each sensor modality under varying operating conditions, accounting for the influence of exogenous variables.
  • First heterogeneous datasets for load virtual sensing: We contribute two new, publicly available datasets, specifically designed to advance research in virtual sensing with multimodal sensor data and varying operating conditions, including: a real-world experimental bearing dataset and a simulated train-bridge-track interaction dataset.
@@ 2 Related Works
In this section, we review state-of-the-art virtual sensing methods, covering both model-based and data-driven approaches, discuss recent advances in deep time series modeling for handling multiscale dynamics and exogenous variables, and explore advancements in graph neural networks for time series modeling and their emerging applications in virtual sensing.
@@
Multiscale modeling and exogenous variables. Virtual sensing must address the challenges of heterogeneous temporal dynamics and exogenous variable influence. Deep time series models, while often not designed as regression models for virtual sensing, have explored multiscale modeling and exogenous variable integration [35]. For instance, Pyraformer [36] utilizes a pyramidal attention module to capture multi-resolution features in time series for forecasting tasks. Similarly, TimeMixer [37], an MLP-based architecture, leverages disentangled multiscale series for past extraction and future prediction. However, these methods have not been applied to virtual sensing problems. Recently, incorporating exogenous variables into time series models has gained traction. For instance, Zhao et al. [38] incorporate exogenous variables as operating condition context to improve model robustness for anomaly detection. TimeXer [39] enhances transformer architectures by incorporating exogenous information through embedding layers, patch-wise self-attention, and variate-wise cross-attention. Similarly, Theiler et al. [40] integrate exogenous variables from future time steps to improve the forecasting accuracy of transformer models. However, these approaches do not explicitly model how exogenous variables can alter the frequency or scale of endogenous variables, a crucial aspect for accurate virtual sensing.
Graph neural networks (GNNs) were originally developed for static graph representations, with applications in fields such as computational chemistry [41] and social networks [42, 43]. Leveraging their ability to capture complex spatial dependencies, GNNs were later adapted to time-series data, giving rise to temporal GNNs. These models have rapidly expanded into the Industrial Internet of Things (IIoT) [44], where effectively model spatial-temporal dependencies in sensor networks is crucial for downstream tasks such as anomaly and fault detection [19, 45]. For example, a GNN-based model leveraging graph attention mechanisms has shown promise in identifying anomalies within multivariate time series [46]. Another approach employed sensor embeddings for graph construction and prediction-based anomaly detection [47]. Recently, a GNN model incorporating dynamic edges has further improved fault detection by capturing evolving relationships between time series, while also considering operating conditions [38]. However, a persistent challenge remains in addressing heterogeneous temporal dynamics often present in sensor data. Beyond anomaly detection, the inherent ability of GNNS to model interactions within sensor networks has shown promise in a board range of applications. These include predicting remaining useful life of bearing [48], forecasting air quality  [49], estimating states in hydro power plants [50], and calibrating ozone sensors using low-cost sensor networks.
@@ 3.2 Framework Overview
Figure 1: Architecture of the proposed Heterogeneous Temporal Graph Neural Network (HTGNN) for virtual sensing. HTGNN effectively infers target variable $\mathbf{y}^{t}$ by capturing complex relationships between heterogeneous sensor nodes with diverse modalities, including low-frequency signals ($\mathbf{X_{L}}^{t_{l}:t}$, blue circles), high-frequency signals ($\mathbf{X_{H}}^{t_{l}:t}$, orange squares) and exogenous variables ($\mathbf{W}^{t_{l}:t}$, dark blue triangle). The model incorporates three key stages: (1) Operating condition aware dynamics modeling: A Multilayer Perceptron (MLP) extracts operating conditions, informing both a Gated Recurrent Unit (GRU) for low-frequency dynamics and a gated Convolutional Neural Network (CNN) for high-frequency dynamics extraction, capturing operating condition-aware temporal patterns within each signal type. (2) Heterogeneous interaction learning: Specialized Graph Neural Networks (GNNs) learn complex inter- and intra-modality interactions between heterogeneous sensor nodes. (3) Target variable inference: A Bidirectional Long Short-Term Memory (BiLSTM) network followed by an MLP analyzes the combined information from the graph embedding to predict the target variable.
@@ 3.5 Heterogeneous Interaction Modelling
To effectively capture the complex relationships between sensor nodes and explicitly account for the influence of operating conditions on these interactions, our HTGNN model strategically models heterogeneous interactions within the temporal graph. We address two distinct types of interactions: intra-modality interactions among nodes with similar signal characteristics (e.g., low-frequency to low-frequency, high-frequency to high-frequency) and inter-modality across nodes with different signal characteristics (e.g., low-frequency to high-frequency). This approach to interaction modeling leverages the node dynamics previously extracted in Sec. 3.4 (low-frequency nodes from Eq. 9, high-frequency nodes from Eq. 14). This method ensures that our model comprehensively integrates both intra- and inter-modality dynamics, enhancing the accuracy and relevance of the sensor data analysis while respecting the signal characteristics.
@@ 4.2 Case Study 2: Bridge
In this case study, we evaluate our proposed HTGNN model for bridge health monitoring, specifically focusing on estimating live load on the bridge deck from displacement and acceleration signals. Live load refers to the dynamic forces exerted on the bridge due to traffic, such as the weight of vehicles passing over it. Accurate live load estimation is essential for assessing structural integrity, identifying potential overloads or fatigue damage, and making informed decisions about maintenance and repair [55]. This capability is essential for ensuring the safety and longevity of bridge structures.
@@ 5 Evaluation Setup
To provide a comprehensive evaluation of our proposed method, we benchmark its performance against several established approaches for virtual sensing [12], each selected for its relevance to the problem domain. We chose Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) as our primary baselines due to their proven effectiveness in handling temporal dynamics [12], a key aspect of our virtual sensing task with heterogeneous temporal dynamics. CNNs are well suited for local dynamics modeling and frequency domain processing, while RNNs are capable of capturing long-term dependencies and overall dynamics modeling. Both these capabilities are crucial for effectively capturing the complex temporal relationships in our problem. Specifically, we compare against the Bidirectional LSTM (BiLSTM) due to its widespread use in sequence modeling, and the 1D Convolutional Neural Network (1D-CNN) for its effectiveness in signal processing tasks. The 1D Gated CNN (1D-GCNN) is included to assess the impact of gating mechanisms similar to those in our approach. To assess the benefits of multiscale modeling, we include TimeMixer [37], originally designed for forecasting tasks. We adapt its architecture for our regression task by removing the decoder component and introducing a regression head, following the implementation provided in [35] adapted for classification tasks. Finally, we include the Multivariate Time-series Graph Attention Network (MTGAT) , originally introduced by [46] as an anomaly detection framework and later adapted by [59] for the sensor calibration task, as a state-of-the-art homogeneous temporal graph regression model. This serves as a benchmark for the effectiveness of our heterogeneous graph-based approach. For each model, we added a head that converts from the final hidden dimension of the model to the output, implemented as a three-layered MLP. All models utilized the SiLU (Sigmoid Linear Unit) activation function. For further details on the hyperparameter tuning of the baseline methods and our proposed HTGNN model, please refer to A.3.
@@ 6.1.1 Overall Quantitative Performance
Baseline method performance. The complex, interdependent interactions between temperature and vibration measurements in bearing load prediction pose a challenge for many models. These interactions vary not only with load but also with rotational speed. While higher loads generally decrease vibration magnitudes [63] but increase temperature change rates [64], the influence of load on these change rates can be altered by the rotational speed. At higher speeds, both temperature and vibration magnitudes often exhibit steeper increases or decreases in response to load changes. Comparison models, such as MTGAT, which assume homogeneous relations between sensor nodes, struggle to capture these heterogeneous dynamics. Similarly, sequence-based models like 1D-CNN and BiLSTM, primarily focusing on temporal dependencies within a single sensor modality, also face limitations in modeling the complex relationships between temperature and vibration across varying rotational speeds. TimeMixer demonstrates greater stability in these worst-case scenarios, with the highest MAPE for both axial and radial load prediction remaining below 25%. This stability highlights the benefits of multiscale modeling in improving model robustness. However, while TimeMixer can capture some multiscale dynamics, it does not explicitly capture inter-scale interactions or fully account for the impact of exogenous variables like rotational speed, which limits its performance under rare operating conditions and leads to worse predictions at lower speeds.
@@
Benefits of heterogeneous interaction modeling. In contrast, HTGNN’s architecture explicitly models these heterogeneous relationships and their dependence on operating conditions. This enables HTGNN to effectively extract information from both temperature and vibration sensors across varying rotational speeds, resulting in improved load prediction. By capturing these complex dependencies, HTGNN outperforms most baseline models across all rotational speeds and load types, as evidenced by the lowest NRMSE values with tight confidence intervals (Tab. 1) and the lowest maximum MAPE values (Tab. 2). The low variance in HTGNN’s performance metrics further underscores its robustness and reliability, making it a promising solution for virtual load sensing in complex industrial systems.
@@ 6.1.2 In-Depth Analysis of Load Predictions
Challenges in load prediction. The right two plots of Fig. 8 illustrate the varying degrees of deviation from true load values observed in the performance of baseline models, particularly for axial loads. The relatively flat frequency spectra observed at low speeds (Figure 7) indicate a lower signal-to-noise ratio, which may hinder the performance of frequency-based models like 1D-CNN and 1D-GCNN that rely on prominent frequency-domain features. While 1D-GCNN’s gating mechanism offers some mitigation compared to 1D-CNN, both models remain susceptible to significant deviations under these conditions. MTGAT, a homogeneous GNN, also exhibits occasional extreme errors. This might be due to its aggregation of inherently heterogeneous features (e.g., temperature and vibration), potentially hindering its ability to extract meaningful representations. Additionally, BiLSTM, as a sequential model, may face challenges in extracting meaningful frequency information from vibration signals, particularly when signal trends are less apparent. TimeMixer, while generally performing well, exhibits sensitivity to changing operating conditions (i.e. during shifts in rotational speed). This sensitivity likely stems from its inability to explicitly account for the impact of exogenous variables on the measured signals.
@@
Worst case performance assessment. To further assess the robustness of each model, we examine their worst-case performance under challenging operating conditions. Fig. 9 presents the average MAPE for their two most challenging scenarios across five runs. The error bars represent the 95% confidence intervals. Notably, these categories often involve low axial and radial loads, which inherently lead to higher MAPE values due to the nature of this metric (i.e., dividing by a small actual value). For axial load prediction, 1D-CNN and 1D-GCNN consistently struggle, particularly under low axial loads and high rotational speeds. These conditions often lead to significant fluctuations and variations in vibration signal magnitudes. CNN-based methods may not capture effectively due to their primary focus on the amplitude of frequency components rather than the overall signal trends and changes, which are crucial for accurate load prediction. While 1D-GCNN’s gating mechanism offers some improvement over 1D-CNN by modulating extracted features based on rotational speed, it does not consistently rectify this limitation across all operating conditions. Similarly, BiLSTM exhibits high axial MAPE under both high radial loads with high speeds and low radial loads with low speeds. These conditions lead to extreme vibration amplitudes (high or low), which BiLSTM, as a sequential model, may struggle to handle effectively due to the complexities in extracting relevant features from such signals. These findings highlight the complex nonlinear relationship between vibration, temperature, rotational speed, and load, posing a particular challenge for baseline models that lack the ability to effectively capture both high-frequency and low-frequency features simultaneously. TimeMixer demonstrates greater stability in these worst-case scenarios, with the highest MAPE for both axial and radial load prediction remaining below 25%. This stability highlights the benefits of multiscale modeling in improving model robustness. However, TimeMixer, like the other baselines, does not explicitly account for the influence of exogenous variables on the measured signals and therefore performs worse under rarely represented operating conditions, such as low load and low rotational speed.
@@ 6.2 Bridge Load Prediction
6.2.1 Overall Performance
Tab. 3 presents the model performance in predicting train loads across different temperature categories using Normalized Root Mean Squared Error (NRMSE) with 95% confidence intervals. Notably, all models exhibit a slight increase in NRMSE as temperature decreases. This could be attributed to the inherent difficulty of modeling bridge behavior at lower temperatures or to the fact that the training dataset was not sufficiently representative and did not contain enough observations of bridge behavior under such conditions. HTGNN consistently demonstrates competitive performance across all temperature ranges, showing particular strength at lower temperatures where other models experience higher NRMSE. 1D-GCNN, with its gating mechanism, also exhibits strong performance, particularly at higher temperatures. 1D-CNN and BiLSTM generally perform well, except at low temperatures. MTGAT and TimeMixer consistently show the highest NRMSE values. This underperformance, in contrast to their competitive results on the bearing dataset, may be attributed to the increased heterogeneity of the bridge sensor dataset. Unlike the bearing dataset, which is dominated by low-frequency temperature signals, the bridge dataset contains a balanced distribution of high-frequency (acceleration) and low-frequency (displacement) signals, further compounded by the pronounced impact of the exogenous variable (temperature) on these signals. The increased heterogeneity could pose challenges for MTGAT’s homogeneous graph structure as well as TimeMixer’s limited ability to capture exogenous influences, limiting their ability to fully capture the complex interactions between different sensor modalities.